OpenStudy (blackops2luvr):

A classroom globe has a diameter of 12 inches. What is the volume of the globe? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (blackops2luvr):

@igreen

OpenStudy (anonymous):

Easy... -.-

OpenStudy (anonymous):

The formula for the volume of a sphere is 4/3 x 3.14 x radius^3

OpenStudy (blackops2luvr):

lol

OpenStudy (anonymous):

Since the diameter is 12, the radius would be half of it, or 6. So can you figure out 4/3 x 3.14 x 6^3?

OpenStudy (blackops2luvr):

904.32?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Good job person who got me banned. :3

OpenStudy (igreen):

Formula: \(\dfrac{4}{3} \pi r^3\) Plug in what we know: \(\dfrac{4}{3} (3.14) (6)^3\) Simplify exponent: \(\dfrac{4}{3} (3.14) (216)\) Multiply: \((4.19)(216)\) Can you multiply that? @blackops2luvr

OpenStudy (blackops2luvr):

XD

OpenStudy (anonymous):

@iGreen He already got the answer. :P

OpenStudy (blackops2luvr):

lol no i wana give you guys both a medal.